# Denoising Functional Maps: Diffusion Models for Shape Correspondence

### v1

@@ Abstract
Estimating correspondences between pairs of deformable shapes remains a challenging problem. Despite substantial progress, existing methods lack broad generalization capabilities and require category-specific training data. To address these limitations, we propose a fundamentally new approach to shape correspondence based on denoising diffusion models. In our method, a diffusion model learns to directly predict the functional map, a low-dimensional representation of a point-wise map between shapes. We use a large dataset of synthetic human meshes for training and employ two steps to reduce the number of functional maps that need to be learned. First, the maps refer to a template rather than shape pairs. Second, the functional map is defined in a basis of eigenvectors of the Laplacian, which is not unique due to sign ambiguity. Therefore, we introduce an unsupervised approach to select a specific basis by correcting the signs of eigenvectors based on surface features. Our approach achieves competitive performance on standard human datasets, meshes with anisotropic connectivity, non-isometric humanoid shapes, as well as animals compared to existing descriptor-based and large-scale shape deformation methods. The source code and the datasets can be found on our project page: https://alekseizhuravlev.github.io/denoising-functional-maps/.
Figure 1: We propose DenoisFM, a novel framework for predicting shape correspondences in the form of functional maps using denoising diffusion models. (Left:) Challenging examples our method can handle, with color-coded correspondences. (Right:) By sequentially denoising samples of random noise, the diffusion model can predict the correct functional map between a pair of shapes.
@@ 1 Introduction
A long-standing problem in computer graphics and vision is to find corresponding points in a pair of 3D shapes [9, 49, 19, 64, 69, 11, 42]. The point map can then be used for texture [14], segmentation [63] and deformation transfer [68], statistical shape analysis [13], and construction of parametric models [37, 55]. A popular approach to solving the correspondence problem involves the functional map framework [49], which represents the correspondence as a small matrix that maps the eigenfunctions of the Laplace operator on each shape. This representation is compact and allows to enforce constraints on the correspondences, such as bijectivity or area preservation. The functional map framework has been extended in numerous works, such as matching shapes with partiality [58, 4], analyzing the properties of shape collections [13] or improving accuracy through iterative refinement [44, 19, 57].
Previous functional map approaches focused on using small datasets, rarely exceeding a hundred unique shapes. This favored training models with lightweight architectures that could learn as much as possible with limited data. At the same time, they would be highly category-specific, e.g. suitable for a particular class of human shapes but lacking generalization to humanoids. To develop a unified model that is applicable to various tasks, we draw attention to denoising diffusion models  [66, 67, 25], which have impressive generalization ability when trained on large amounts of data. In this paper, we apply this approach to correspondence estimation using functional maps.
Based on the idea of sequentially denoising images blurred with Gaussian noise, diffusion models have achieved unprecedented performance in high-fidelity, diverse, and controllable image synthesis [53]. The scalable U-Net architecture [60] allows them to be trained on large datasets with hundreds of thousands of high-resolution images. The application of diffusion models to shape matching has been limited so far, mainly focusing on feature extraction from pre-trained foundational models [17, 1]. Interestingly, functional maps appear well-suited to bridge diffusion models and shape correspondence. They are small-sized single-channel matrices, making standard architectures developed for images readily applicable to them. Large amounts of shape data for training are then available through parametric models such as SMPL [37], an approach that has already been used for matching tasks [23, 72] and is also suitable for diffusion models.
@@
An important property of functional maps is that they are defined in the basis of the first $n$ eigenvectors of the Laplacian operator for each shape. The basis is not unique due to the sign ambiguity of the eigenvectors, and there are potentially $2^{n}$ possible basis combinations for each shape. To address this problem, we propose an unsupervised approach which selects the signs of the basis vectors using per-vertex features extracted in an unsupervised way (see Sec. 3.2). This reduces the number of possible functional maps between shapes, further lowering the training complexity.
@@
Our model performs on par with existing shape matching methods on human benchmark datasets, achieves generalization to non-isometric humanoid shapes without learning on them, and is also applicable to animal shapes. See our project page for the source code and the datasets of functional maps used to train our models.
@@ 2 Related Work
Since for any eigenvector $\phi_{i}$, its negative version $-\phi_{i}$ is also an eigenvector, the Laplace-Beltrami eigenbasis suffers from sign ambiguity. Additionally, eigenvalues can have higher multiplicity and make the intrinsic ordering non-unique. This symmetry needs to be taken into account when processing the eigenvectors [63, 48, 8].
@@ 3 Background
Given a shape $S$ represented as a triangular mesh with $v$ vertices, we compute its Laplace matrix $L$ [45] and the first $n$ eigenvectors of $L$ corresponding to the eigenvalues $\lambda_{1}\leqslant\ldots\leqslant\lambda_{n}$. $A$ is then the diagonal vertex-area matrix associated with $L$, computed by assigning one-third of the total area of all faces incident to the corresponding vertex.
@@ 3.2 Sign Ambiguity
Sign ambiguity is a special case of basis ambiguity when all eigenvalues are distinct. The more general basis ambiguity occurs if an eigenvalue $\lambda_{i}$ has multiplicity degree $d_{i}>1$. Then any $d_{i}$ vectors forming an orthogonal basis in the subspace spanned by its eigenvectors $\left(\phi_{i+1},\phi_{i+2}\ldots\phi_{i+d_{i}}\right)$ are also valid eigenvectors. Although the basis ambiguity is difficult to deal with, in practice high multiplicities do not occur often and are less common in low-order eigenvectors [29, 50].
@@ 4 Denoising Functional Maps
(a) Functional map prediction using a diffusion model. For each unique shape in the dataset, we obtain surface features using a feature extractor, as well as a sign-corrected eigenbasis (Sec. 4.2). Both are used for conditional prediction of the template-wise functional map using a diffusion model, performed separately for each shape. The pairwise functional maps are obtained through the map composition property, upsampled using ZoomOut [44], and converted to pointwise maps.
(b) Learned sign correction. The eigenvectors returned by a numerical eigensolver have random signs. To select a specific sign, we project the eigenvectors onto the learned correction vector obtained with a feature extractor, and make the projection positive. The sign correction network is trained in an unsupervised manner: learning to correct eigenvectors that have random signs.
Figure 2: Overview of DenoisFM. (a) A pair of shapes is matched using a DDPM model. (b) The signs of eigenvectors are corrected through a learned projection vector.
@@ 4.1 Denoising Diffusion Component
4.1.1 Overview
We next describe the training and inference of the diffusion component of DenoisFM. For training, we use a large collection of shapes and a template. For each shape $S$ in the dataset, we compute an eigenbasis $\Phi$, a template-wise functional map $C_{1T}$, and conditioning $y=y\left(S,\Phi\right)$ which incorporates geometric details of the input shape and information about the basis. The exact form of conditioning will be specified in Sec. 4.3. We use the training objective of denoising diffusion probabilistic models (DDPMs) [25]:
$\mathcal{L}_{\text{DDPM}}=E_{x,\epsilon\sim N(0,1),t}\left[\left\|\epsilon-\epsilon_{\theta}\left(x_{t},t,y\right)\right\|_{2}^{2}\right],$ (3)
where $\epsilon_{\theta}$ is the amount of noise predicted by a 2D UNet [60], $x_{t}$ is a noisy version of the input $x_{0}\coloneq C_{1T}$, and $t$ is the time step uniformly sampled from $\{1,\ldots,T\}$.
The inference pipeline consists of two stages, see Fig. 2(a). In the template stage, for each individual shape in the evaluation dataset, we obtain an eigenbasis, calculate the conditioning, and predict the template-wise functional map using the diffusion model. Note that the inference of DDPMs is time-consuming due to the large number of denoising steps and can take several hours for high-dimensional functional maps. Because the predicted maps refer to the template, we need to run the denoising process only once for each shape.
@@ 4.1.2 Post-Processing of the Samples
Diffusion models are inherently probabilistic and their outputs may differ given the same conditioning but different initial noise [25]. To address this, we repeat the denoising process multiple times and filter out the low-quality results. The selection is based on the Dirichlet energy of the point-wise map [40, 21]: an accurate map should be smooth, with neighboring vertices on one shape matching neighboring vertices on the other. Therefore, we repeat both the template step and the pairwise step to produce multiple point-wise maps. We can then either report the map with the lowest Dirichlet energy or combine the results of several smoothest maps. To do this, for each query point on the source mesh, we gather potential matching points on the target mesh and select the one with the smallest total distance to the other candidates (see Supplement B.3). We found the map selection step to be very important in determining the final accuracy of the method.
@@ 4.2 Learned Sign Correction
4.2.1 Intuition
The diffusion model should be conditioned on the basis eigenvectors and predict the functional map for that particular basis. Theoretically, the model could be trained using random basis combinations, as was done in Dwivedi et al. [18] for graph networks. However, this would dramatically increase the amount of training data.
Hence, we follow a more effective approach: we select a specific basis by fixing the sign of the Laplacian eigenvectors on each shape. Previous works infer the sign from the underlying data using the projection operator (see Sec. 3.2). Therefore, suppose that $\Sigma\in\mathbb{R}^{v\times n}$ is a set of per-vertex features inferred from the mesh geometry, one for each eigenvector. The sign of the eigenvectors can be selected by projecting the features $\Sigma$ onto the eigenvectors $\Phi$ and making the corresponding projections positive. This can be written as:
@@ 4.2.1 Intuition
where “$\operatorname{diag}(\cdot)$” and “$\operatorname{sign}(\cdot)$” denote matrix diagonal and sign extraction operators, $A$ is the diagonal vertex-area matrix. To illustrate this, suppose we are given a human mesh and a single feature vector $\varsigma_{i}$, which is a binary vector with ones for points on the left hand, and zeros otherwise. After performing the procedure described in Eq. (4), the corrected eigenvector $\hat{\phi}_{i}$ will be predominantly positive on the left hand.
@@ 4.2.2 Training the Sign Corrector
In practice, the signs of eigenvectors should be selected using priors learned by a neural network. We define the learned correction vector $\Sigma:=\Theta\left(S\right)$, where $\Theta$ is a trainable per-vertex feature extractor. If we perform the eigendecomposition with a numerical solver twice on the same mesh and get two sets of eigenvectors $\Phi_{1}$ and $\Phi_{2}$, due to the sign ambiguity, the difference between them can be written as $\Phi_{2}=\Phi_{1}\,\sigma$, where $\sigma\in\{-1,1\}^{n}$ is the ground-truth sign difference. Now we perform the sign correction using Eq. (4), after which we expect the corrected eigenvectors to be equal to each other:
@@
where $\hat{\sigma}_{1},\hat{\sigma}_{2}\in\{-1,1\}^{n}$ are the predicted combinations of correction signs. After substitution and using $\hat{\sigma}_{1}\,\hat{\sigma}_{1}=I$, we obtain:
@@
Note that the $\operatorname{sign}$ operator in Eq. (4) is not differentiable. Therefore, during training, we set $\hat{\sigma}_{1}$ and $\hat{\sigma}_{2}$ to be the diagonal elements of the projections, rather than their signs, assuming that the correction vectors and eigenvectors have unit norm. The training objective of the feature extractor (Eq. (7)) then forces the elements of $\hat{\sigma}_{1}$ and $\hat{\sigma}_{2}$ to be as close to $-1$ or $1$ as possible. The absolute value can be interpreted as confidence in the predicted sign: values close to ±$\pm 1$ mean high confidence.
In practice, we found that a network trained this way converges quite quickly on a dataset of human meshes. Our method returns the same corrected eigenbasis 95–99% of the time for the first hundred Laplacian eigenvectors. The accuracy of the sign correction and its impact on the distribution of functional maps are discussed in Supplement A.
@@ 4.3 Choice of Conditioning
Conditioning of the diffusion model should include two parts: the basis information and the geometric features of the shape. The former can be provided simply via the sign-corrected eigenvectors $\hat{\Phi}$. They additionally represent the spectral embedding of the shape and convey the full intrinsic information. As for the latter, during the training of the sign-correction network, the useful features will be encoded in the learned correction vector $\Sigma$. Therefore, we propose to project the learned correction vectors onto the sign-corrected eigenvectors:
@@ 5 Training and Experimental Results
This section describes the steps for generating the training data for our model (Sec. 5.1) as well as the experimental results on several challenging datasets (Secs. 5.25.5). The evaluation of the sign correction network, additional comparison, and the ablation studies can be found in Supplement A and C. Since our model is designed to be trained on large and diverse shape collections, it can be applied to multiple (reminiscent) shape classes without re-training. This reduces the need for small category-specific training sets. We focus on datasets with human-like shapes as their combined volume allows versatile training and comparison of our new DDPM-based approach with the previous state of the art.
@@ 5.1 Training Data and Implementation
Our model consists of two independent components: a denoising diffusion model that predicts the functional map and a sign correction network based on a per-vertex feature extractor. To reduce computational overhead, we train them separately on the respective datasets.

### v2

@@ Abstract
Estimating correspondences between pairs of deformable shapes remains a challenging problem. Despite substantial progress, existing methods lack broad generalization capabilities and require category-specific training data. To address these limitations, we propose a fundamentally new approach to shape correspondence based on denoising diffusion models. In our method, a diffusion model learns to directly predict the functional map, a low-dimensional representation of a point-wise map between shapes. We use a large dataset of synthetic human meshes for training and employ two steps to reduce the number of functional maps that need to be learned. First, the maps refer to a template rather than shape pairs. Second, the functional map is defined in a basis of eigenvectors of the Laplacian, which is not unique due to sign ambiguity. Therefore, we introduce an unsupervised approach to select a specific basis by correcting the signs of eigenvectors based on surface features. Our model achieves competitive performance on standard human datasets, meshes with anisotropic connectivity, non-isometric humanoid shapes, as well as animals compared to existing descriptor-based and large-scale shape deformation methods. See our project page [^1] Project page: https://alekseizhuravlev.github.io/denoising-functional-maps/ for the source code [^2] Source code: https://github.com/alekseizhuravlev/denoising-functional-maps/ and the datasets.
Figure 1: We propose DenoisFM, a novel method for predicting shape correspondences in the form of functional maps using denoising diffusion models. (Left:) Challenging examples our method can handle, with color-coded correspondences. (Right:) By sequentially denoising samples of random noise, the diffusion model can predict the correct functional map between a pair of shapes.
@@ 1 Introduction
A long-standing problem in computer graphics and vision is to find corresponding points in a pair of 3D shapes [9, 49, 19, 64, 69, 11, 42]. The point map can then be used for texture [14], segmentation [63] and deformation transfer [68], statistical shape analysis [13], and construction of parametric models [37, 55]. A popular approach to solving the correspondence problem involves the functional map framework [49], which represents the correspondence as a small matrix that maps between the eigenfunctions of the Laplace operator on each shape. This representation is compact and allows to regularize constraints on the correspondences, such as bijectivity or area preservation. The functional map framework has been extended in numerous works, such as matching shapes with partiality [58, 4], analyzing the properties of shape collections [13] or improving accuracy through iterative refinement [44, 19, 57].
Previous functional map approaches focused on using small datasets, rarely exceeding a hundred unique shapes. This favored training models with lightweight architectures that could learn as much as possible with limited data, but these methods may face challenges with generalization to broad shape categories. To develop a unified model that is applicable to various tasks, we draw attention to denoising diffusion models  [66, 67, 25], which have impressive generalization ability when trained on large amounts of data. In our work, we apply this approach to correspondence estimation using functional maps.
Based on the idea of sequentially denoising samples of random noise into samples from the data distribution, diffusion models have achieved unprecedented performance in high-fidelity, diverse, and controllable image synthesis [53]. The scalable U-Net architecture [60] allows them to be trained on large datasets with hundreds of thousands of high-resolution images. The application of diffusion models to shape matching has been limited so far, mainly focusing on feature extraction from pre-trained foundational models [17, 1]. Interestingly, functional maps appear well-suited to bridge diffusion models and shape correspondence. They are small-sized single-channel matrices, making standard architectures developed for images readily applicable to them. Large amounts of shape data for training are then available through parametric models such as SMPL [37], an approach that has already been used for matching tasks [23, 72] and is also suitable for diffusion models.
@@
An important property of functional maps is that they are defined in the basis of the first $n$ eigenvectors of the Laplacian operator for each shape. The basis is not unique due to the sign ambiguity of the eigenvectors, and there are potentially $2^{n}$ possible basis combinations for each shape. To address this problem, we propose an unsupervised approach which selects the signs of the basis vectors using per-vertex features extracted in an unsupervised way (Sec. 3.2). This reduces the number of possible functional maps between shapes, further lowering the training complexity.
@@
Our model performs on par with existing shape matching methods on human benchmark datasets, achieves generalization to non-isometric humanoid shapes without learning on them, and is also applicable to animal shapes. See our project page1 for the source code2 and the datasets of functional maps used to train our models.
@@ 2 Related Work
Since for any eigenvector $\phi_{i}$, its negative version $-\phi_{i}$ is also an eigenvector, the Laplace-Beltrami eigenbasis has an inherent sign ambiguity. Additionally, the eigenvalues can have higher multiplicity and make the intrinsic ordering non-unique. This symmetry needs to be taken into account when processing the eigenvectors [63, 48, 8].
@@ 3 Background
Given a shape $S$ represented as a triangular mesh with $v$ vertices, we compute its Laplace matrix $L$ [45] and the first $n$ eigenvectors of $L$ corresponding to the eigenvalues $\lambda_{1}\leqslant\ldots\leqslant\lambda_{n}$. $A$ is the diagonal vertex-area matrix associated with $L$, computed by assigning one-third of the total area of all faces incident to the corresponding vertex.
@@ 3.2 Sign Ambiguity
Sign ambiguity is a special case of basis ambiguity when all eigenvalues are distinct. The more general basis ambiguity occurs if an eigenvalue $\lambda_{i}$ has multiplicity degree $d_{i}>1$. Then any $d_{i}$ vectors forming an orthogonal basis in the subspace spanned by its eigenvectors $\left(\phi_{i+1},\phi_{i+2}\ldots\phi_{i+d_{i}}\right)$ are also valid eigenvectors. Although the basis ambiguity is difficult to deal with, in practice high multiplicities do not occur often in low-order eigenvectors [29, 50].
@@ 4 Denoising Functional Maps
(a) Functional map prediction using a diffusion model. For each unique shape $S$ in the dataset, we obtain feature vectors $\Sigma$ using a feature extractor $\Theta$, as well as a sign-corrected eigenbasis $\hat{\Phi}$ (Fig. 2(b), Sec. 4.2). Both are used for conditional prediction of the template-wise functional map $C_{1T}$ using a diffusion model, performed separately for each shape. The pairwise functional maps $C_{12}$ are obtained through the map composition property, upsampled using ZoomOut [44], and converted to pointwise maps $\Pi_{21}$.
(b) Learned sign correction. The eigenvectors $\Phi$ returned by a numerical eigensolver have random signs. To select a specific sign, we project the eigenvectors onto the feature vectors $\Sigma$ obtained with a feature extractor $\Theta$, and make the projection positive. The sign correction network is trained in an unsupervised manner: learning to correct eigenvectors that have random signs. The learned features $\Sigma$ are also used for the conditional prediction of the template-wise functional map using a diffusion model (Fig. 2(a)).
Figure 2: Overview of DenoisFM. (a) A pair of shapes is matched using a DDPM model. (b) The signs of eigenvectors are corrected through learned features.
@@ 4.1 Denoising Diffusion Component
We next describe the training and inference of the diffusion component of DenoisFM. For training, we use a large collection of shapes and a template. For each shape $S$ in the dataset, we compute an eigenbasis $\Phi$, a template-wise functional map $C_{1T}$, and conditioning $y=y\left(S,\Phi\right)$ which incorporates geometric details of the input shape and information about the basis. The exact form of conditioning will be specified in Sec. 4.3. We use the training objective of denoising diffusion probabilistic models (DDPMs) [25]: $\mathcal{L}_{\text{DDPM}}=E_{x,\epsilon\sim N(0,1),t}\left[\left\|\epsilon-\epsilon_{\theta}\left(x_{t},t,y\right)\right\|_{2}^{2}\right],$ where $\epsilon_{\theta}$ is the amount of noise predicted by a 2D UNet [60], $x_{t}$ is a noisy version of the input $x_{0}\coloneq C_{1T}$, and $t$ is the time step uniformly sampled from $\{1,\ldots,T\}$.
The inference pipeline consists of two stages, see Fig. 2(a). In the template stage, for each individual shape $S$ in the evaluation dataset, we obtain an eigenbasis $\Phi$, calculate the conditioning $y$, and predict the template-wise functional map $C_{1T}$ using the diffusion model. Note that the inference of DDPMs is time-consuming due to the large number of denoising steps and can take several hours for high-dimensional functional maps. Because the predicted maps refer to the template, we need to run the denoising process only once for each shape.
@@ 4.1.1 Post-Processing of the Samples
Diffusion models are inherently probabilistic and their outputs may differ given the same conditioning but different initial noise [25]. To address this, we repeat the denoising process multiple times and filter out the low-quality results. The selection is based on the Dirichlet energy of the point-wise map [40, 21]: an accurate map should be smooth, with neighboring vertices on one shape matching neighboring vertices on the other. Therefore, we repeat both the template step and the pairwise step to produce multiple point-wise maps. We can then either report the map with the lowest Dirichlet energy or combine the results of several smoothest maps. To do this, for each query point on the source mesh, we gather potential matching points on the target mesh and select the one with the smallest total distance to the other candidates (see Supp. B.3). The map selection step has a major impact on the accuracy of our method.
@@ 4.2 Learned Sign Correction
The diffusion model needs to be conditioned on the basis eigenvectors and predict the functional map for that particular basis. Theoretically, the model could be trained using random basis combinations, as was done in Dwivedi et al. [18] for graph networks. However, this would significantly increase the amount of training data.
Hence, we follow a more effective approach: we select a specific basis by fixing the sign of the Laplacian eigenvectors on each shape. Previous works infer the sign from the underlying data using the projection operator (see Sec. 3.2). Therefore, suppose that $\Sigma:=\Theta\left(S\right)\in\mathbb{R}^{v\times n}$ is a set of per-vertex feature vectors inferred from the geometry of shape $S$ by a trainable feature extractor $\Theta$, one for each eigenvector. The sign of the eigenvectors can be selected by projecting the features $\Sigma$ onto the eigenvectors $\Phi$ and making the corresponding projections positive:
@@
where “$\operatorname{diag}(\cdot)$” and “$\operatorname{sign}(\cdot)$” denote matrix diagonal and sign extraction operators, $A$ is the diagonal vertex-area matrix. To illustrate, suppose we have a human mesh and a single feature vector $\varsigma_{i}$, which is a binary vector with ones for points on the left hand, and zeros otherwise. After performing the procedure described in Eq. (3), the corrected eigenvector $\hat{\phi}_{i}$ will be mostly positive on the left hand.
@@ 4.2.1 Training the Sign Corrector
The sign correction network is trained in an unsupervised manner by learning to correct eigenvectors that have random signs. If we perform the eigendecomposition with a numerical solver twice on the same mesh and get two sets of eigenvectors $\Phi_{1}$ and $\Phi_{2}$, due to the sign ambiguity, the difference between them can be written as $\Phi_{2}=\Phi_{1}\,\sigma$, where $\sigma\in\{-1,1\}^{n}$ is the ground-truth sign difference. Now we perform the sign correction using Eq. (3), after which we expect the corrected eigenvectors to be equal to each other:
@@
where $\hat{\sigma}_{1},\hat{\sigma}_{2}\in\{-1,1\}^{n}$ are the predicted combinations of correction signs. After substitution and using $\hat{\sigma}_{1}\,\hat{\sigma}_{1}=I$:
@@
Note that the $\operatorname{sign}$ operator in Eq. (3) is not differentiable. Therefore, during training, we set $\hat{\sigma}_{1}$ and $\hat{\sigma}_{2}$ to be the diagonal elements of the projections, rather than their signs, assuming that the feature vectors and eigenvectors have unit norm. The training objective of the feature extractor (Eq. 6) then forces the elements of $\hat{\sigma}_{1}$ and $\hat{\sigma}_{2}$ to be as close to $-1$ or $1$ as possible. The absolute value can be interpreted as confidence in the predicted sign: values close to ±$\pm 1$ mean high confidence.
In practice, we found that a network trained this way quickly converges on a dataset of human or animal meshes. Our method returns the same corrected eigenbasis 95–99% of the time for the first hundred Laplacian eigenvectors. The accuracy of the sign correction and its impact on the distribution of functional maps are discussed in Supp. A.
@@ 4.3 Choice of Conditioning
Conditioning of the diffusion model should include two parts: the basis information and the geometric properties of the shape. The former can be provided simply via the sign-corrected eigenvectors $\hat{\Phi}$. They additionally represent the spectral embedding of the shape and convey the full intrinsic information. As for the latter, during the training of the sign-correction network, the useful properties will be encoded in the learned feature vectors $\Sigma$. Therefore, we propose to project the feature vectors onto the sign-corrected eigenvectors:
@@ 5 Training and Experimental Results
This section describes the steps for generating the training data for our model (Sec. 5.1) as well as the experimental results on several challenging datasets (Secs. 5.25.5). The evaluation of the sign correction network, additional comparison, implementation details, and the ablation studies can be found in Supp. A to C. Since our model is designed to be trained on large and diverse shape collections, it can be applied to multiple (reminiscent) shape classes without re-training. This reduces the need for small category-specific training sets. We focus on datasets with human-like shapes as their combined volume allows versatile comparison of our approach with the previous works.
@@ 5.1 Training Data and Implementation
Our model consists of two independent components: a denoising diffusion model that predicts the functional map and a sign correction network based on a per-vertex feature extractor, trained separately on their respective datasets.